OpenStudy (anonymous):

SLOPE: I need help, please explain. -8y=7x+4 I don't understand, because I've only done problems without a number in front of the y.

OpenStudy (johnweldon1993):

so that means...get the y by itself...divide this whole equation by -8 :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

u want slope??

OpenStudy (anonymous):

\[\huge \frac{ 7 }{ 8 }\]

OpenStudy (johnweldon1993):

because -8y = 7x + 4 --- --- --- -8 -8 -8 makes y be just y y = -7/8x - 4/8 simplify y = -7/8x - 1/2 :)

OpenStudy (anonymous):

thank you for explaining it to me @johnweldon1993 <3

OpenStudy (anonymous):

\[\huge \frac{ -7 }{ 8 }\]

OpenStudy (johnweldon1993):

@deena123 remember that your dividing by -8

OpenStudy (johnweldon1993):

and yeah not a problem @airrbae glad to help :) feel free to message me or fan me if you need any more help :)